OpenStudy (anonymous):

how do you fina a positive angle less than 2 pi that is coterminal with 16pi/5

OpenStudy (mertsj):

\[\frac{16\pi}{5}=3 \frac{1}{5} \pi -2\pi =1 \frac{1}{5}\pi \]

OpenStudy (mertsj):

And that is the angle that is coterminal with the given angle.